# The $\Lambda_{b}\to\Lambda$ transition form factors in perturbative QCD approach

### v1

@@ III Light-cone distribution amplitudes
  • twist-3
    $V_{1}(x_{i})$ $=$ $0\,,\hskip 128.0374ptA_{1}(x_{i})=-120x_{1}x_{2}x_{3}\phi_{3}^{0}\,$,
    $T_{1}(x_{i})$ $=$ $0\,.$ (28)
  • twist-4
    $S_{1}(x_{i})$ $=$ $6x_{3}(1-x_{3})(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{1}(x_{i})=6x_{3}(1-x_{3})(\xi_{4}^{0}-\xi_{4}^{{}^{\prime}0})\,$,
    $V_{2}(x_{i})$ $=$ $0\,,\hskip 142.26378ptA_{2}(x_{i})=-24x_{1}x_{2}\phi_{4}^{0}\,$,
    $V_{3}(x_{i})$ $=$ $12(x_{1}-x_{2})x_{3}\psi_{4}^{0}\,,\hskip 65.44142ptA_{3}(x_{i})=-12x_{3}(1-x_{3})\psi_{4}^{0}\,$,
    $T_{2}(x_{i})$ $=$ $0\,,\hskip 142.26378ptT_{3}(x_{i})=6(x_{2}-x_{1})x_{3}(-\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,$,
    $T_{7}(x_{i})$ $=$ $-6(x_{1}-x_{2})x_{3}(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,.$ (29)
  • twist-5
    $S_{2}(x_{i})$ $=$ $\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{2}(x_{i})=\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,$,
    $V_{4}(x_{i})$ $=$ $3(x_{2}-x_{1})\psi_{5}^{0}\,,\hskip 79.6678ptA_{4}(x_{i})=-3(1-x_{3})\psi_{5}^{0}\,$,
    $V_{5}(x_{i})$ $=$ $0\,,\hskip 139.41832ptA_{5}(x_{i})=-6x_{3}\phi_{5}^{0}\,$,
    $T_{4}(x_{i})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 36.98866ptT_{5}(x_{i})=0\,$,
    $T_{8}(x_{i})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,.$ (30)
  • twist-6
    $V_{6}(x_{i})$ $=$ $0\,,\hskip 128.0374ptA_{6}(x_{i})=-2\phi_{6}^{0}\,$,
    $T_{6}(x_{i})$ $=$ $0\,.$ (31)
The values $\phi_{i}^{0},\psi_{i}^{0},\xi_{i}^{0}$ and $\xi_{i}^{\prime 0}$ in Eqs. (III-III) connect to four independent parameters as follows,
$\phi_{3}^{0}$ $=$ $\phi_{6}^{0}=-f_{\Lambda},\hskip 79.6678pt\phi_{4}^{0}=\phi_{5}^{0}=-\frac{1}{2}(f_{\Lambda}+\lambda_{1}),$
$\psi_{4}^{0}$ $=$ $\psi_{5}^{0}=\frac{1}{2}(f_{\Lambda}-\lambda_{1}),\hskip 45.52458pt\xi_{4}^{0}=\xi_{5}^{0}=\lambda_{2}+\lambda_{3},$
$\xi_{4}^{{}^{\prime}0}$ $=$ $\xi_{5}^{{}^{\prime}0}=\lambda_{3}-\lambda_{2}.$ (32)
Their numerical values and uncertainties have been derived in QCD sum rules Liu:2009mb ,
$f_{\Lambda}$ $=$ $(6.0\pm 0.3)\times 10^{-3}\;\mbox{GeV}^{2},\hskip 71.13188pt\lambda_{1}=(1.0\pm 0.3)\times 10^{-2}\;\mbox{GeV}^{2},$
$|\lambda_{2}|$ $=$ $(0.83\pm 0.05)\times 10^{-2}\;\mbox{GeV}^{2},\hskip 56.9055pt|\lambda_{3}|=(0.83\pm 0.05)\times 10^{-2}\;\mbox{GeV}^{2}.$ (33)
It is noting that the parameters $\lambda_{2}$ and $\lambda_{3}$ in the above parameters are only specifies their absolute values. In the following calculation, we adopt four scheme depending on different choices of the signs of these two parameters.
  • Scheme-I: $\lambda_{2}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-II: $\lambda_{2}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-III: $\lambda_{2}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-IV: $\lambda_{2}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
@@
The PQCD factorization formula for each feynman diagrams are collected in Appendix. Appendix A:Factorization formulas. We present the numerical results of the form factors at $q^{2}=0\>\text{GeV}^{2}$ under four different schemes in Table 1, comparing them with lattice QCD calculations Detmold:2016pkz . The uncertainties in our calculated results originate from the shape parameter in the exponential model LCDAs of $\Lambda_{b}$ baryon and parameters in the $\Lambda$ baryon LCDAs, The more accurate parameters in the baryon LCDAs will help us to further improve the precision of our theoretical predictions. The results in Table 1 show that Scheme-I and Scheme-II are in good agreement with the lattice QCD predictions using the nominal form factor parameters. In particular, Scheme-I also exhibits excellent consistency with the lattice results obtained using the higher-order parameterization. In contrast, Scheme-III and Scheme-IV yield systematically smaller form factors. This implies that the parameter $\lambda_{2}$ plays an important role in correctly reproducing the magnitude of $\Lambda_{b}\to\Lambda$ form factors, whereas $\lambda_{3}$ has less impact. For simplicity, we therefore adopt Scheme-I in the following calculations and analyses.
Table 1: Form factors of $\Lambda_{b}\to\Lambda$ obtained via four schemes, compared with lattice QCD results Detmold:2016pkz , where [1] and [2] represent lattice results with nominal and higher-order parameter choices, respectively.
Scheme-I Scheme-II Scheme-III Scheme-IV lattice QCD[1] Detmold:2016pkz lattice QCD[2] Detmold:2016pkz
$f_{+}(0)$ $0.192\pm 0.058$ $0.130\pm 0.043$ $114\pm 0.040 0.0$$049\pm 0.030$ $0.160\pm 0.024$ $0.202\pm 0.040$
$f_{\perp}(0)$ $0.239\pm 0.072$ $0.161\pm 0.053$ $142\pm 0.050 0.0$$061\pm 0.036$ $0.207\pm 0.037$ $0.241\pm 0.065$
$f_{0}(0)$ $0.192\pm 0.058$ $0.129\pm 0.043$ $114\pm 0.040 0.0$$049\pm 0.029$ $0.156\pm 0.035$ $0.200\pm 0.041$
$g_{+}(0)$ $0.194\pm 0.059$ $0.132\pm 0.044$ $117\pm 0.041 0.0$$053\pm 0.030$ $0.112\pm 0.028$ $0.197\pm 0.036$
$g_{\perp}(0)$ $0.165\pm 0.050$ $0.113\pm 0.038$ $100\pm 0.036 0.0$$46\pm 0.026 0.0$$095\pm 0.035$ $0.169\pm 0.055$
$g_{0}(0)$ $0.195\pm 0.060$ $0.132\pm 0.044$ $117\pm 0.041 0.0$$053\pm 0.030$ $0.166\pm 0.023$ $0.199\pm 0.038$
$h_{+}(0)$ $0.239\pm 0.072$ $0.161\pm 0.053$ $142\pm 0.050 0.0$$061\pm 0.035$ $0.236\pm 0.041$ $0.228\pm 0.072$
$h_{\perp}(0)$ $0.193\pm 0.059$ $0.131\pm 0.044$ $115\pm 0.041 0.0$$051\pm 0.030$ $0.166\pm 0.022$ $0.219\pm 0.043$
$\widetilde{h}_{+}(0)$ $0.165\pm 0.051$ $113\pm 0.038 0.0$$99\pm 0.036 0.0$$045\pm 0.027$ $0.163\pm 0.029$ $0.263\pm 0.050$
$\widetilde{h}_{\perp}(0)$ $0.193\pm 0.059$ $0.131\pm 0.044$ $116\pm 0.041 0.0$$051\pm 0.030$ $0.156\pm 0.027$ $0.235\pm 0.044$
Table 1 presents our calculated helicity form factors, which satisfy the expected endpoint relations $f_{0}(0)=f_{+}(0)$ and $g_{0}(0)=g_{+}(0)$. For comparison with other theoretical approaches, we convert these to Weinberg form factors in Table 2. These results are consistent with the heavy quark effective theory predictions, where the Weinberg form factors for heavy-to-light baryon transitions obey:
$f_{1}^{V}(0)$ $\approx f_{1}^{A}(0)\gg f_{2,3}^{V,A}(0),$
$f_{2}^{TV}(0)$ $\approx f_{2}^{TA}(0)\gg f_{1}^{TV,TA}(0).$ (34)
as clearly demonstrated in Table 2. Within theoretical uncertainties, our form factor calculations agree well with results from other approaches.
Table 2: Theoretical predictions for the form factors at $q^{2}=0$ of the $\Lambda_{b}\rightarrow\Lambda$ transition, where the symbol $\cdots$ denotes that no relevant result has been reported in the literature.
@@
this work $192\pm 0.058 0.0$$39\pm 0.011 0.0$$035\pm 0.010$ $195\pm 0.060 0.0$$37\pm 0.011 0.0$$037\pm 0.011$
Lattice QCD[1]Detmold:2016pkz $0.158$ $0.041$ 0.017-0.017 $0.137$ $0.055$ 0.117-0.117
Lattice QCD[2]Detmold:2016pkz $0.201$ $0.033$ 0.039-0.039 $0.198$ $0.036$ 0.166-0.166
@@
this work $0.038\pm 0.011$ $193\pm 0.059 0.0$$036\pm 0.010$ $0.193\pm 0.059$
Lattice QCD[1]Detmold:2016pkz 0.059-0.059 0.166-0.166 $0.009$ 0.156-0.156
Lattice QCD[2]Detmold:2016pkz 0.008-0.008 0.219-0.219 $0.034$ 0.235-0.235
LCSRWang:2015ndk 0.025-0.025 0.18-0.18 $0.037$ 0.18-0.18
@@
Table 3 presents the numerical results of the form factors $f_{1}^{V,A}(0)$ decomposed by individual Feynman diagrams. Our analysis reveals that diagrams $d_{7}$, $d_{8}$, $d_{13}$, and $d_{14}$ provide the dominant contributions to $f_{1}^{V,A}(0)$.
@@
$f_{1}^{V}(0)$ $1.7\times 10^{-3}$ $1.7\times 10^{-3}$ $2.6\times 10^{-3}$ $1.1\times 10^{-4}$ $9.5\times 10^{-4}$ $7.2\times 10^{-4}$ 0.051 0.044
$f_{1}^{A}(0)$ $1.8\times 10^{-3}$ $1.8\times 10^{-3}$ $2.8\times 10^{-3}$ $1.6\times 10^{-4}$ $1.2\times 10^{-3}$ $1.7\times 10^{-3}$ 0.051 0.045
@@
$f_{1}^{V}(0)$ $1.1\times 10^{-3}$ $9.6\times 10^{-4}$ $5.3\times 10^{-3}$ $1.7\times 10^{-3}$ 0.047 0.053 0.192
$f_{1}^{A}(0)$ $1.1\times 10^{-3}$ $1.0\times 10^{-3}$ $5.4\times 10^{-3}$ $3.3\times 10^{-4}$ 0.048 0.055 0.195
We analyze the contributions of each twist-term combination for these four diagrams in Table 4. The diagrams $d_{7}$ and $d_{8}$ receive their primary contributions from the combination of $\Lambda_{b}$ twist-4 and $\Lambda$ twist-5 terms, and the diagrams $d_{13}$ and $d_{14}$ are predominantly governed by $\Lambda_{b}$ twist-3 and $\Lambda$ twist-4 combinations. These findings demonstrate the essential role of high-power effects in baryon transition form factors at the scale of the b-quark mass.
@@
$\Lambda_{b}$ twist-4 0.0078 $\sim 0$ 0.0427 $\sim 0$ 0.0690 $\sim 0$ 0.0372 $\sim 0$
$d_{13}$ $d_{14}$
$\Lambda_{b}$ twist-2 $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-3 $\sim 0$ 0.0308 $\sim 0$ $\sim 0$ $\sim 0$ 0.0351 $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-4 $\sim 0$ -0.0006 0.0164 $\sim 0$ $\sim 0$ -0.0007 0.0185 $\sim 0$
Recalling the form factors are expressed as the convolution of the perturbatively calculated hard kernels and the non-perturbative baryon LCDAs. We list the numerators of the hard kernels for the diagram $d_{7}$ contributing to the form factor $f_{1}^{V}(0)$ in Table 5, classified by different twist combination. $i=2,3^{+-},3^{-+},4$ stands for the $\Lambda_{b}$ baryon LCDA and $j=3,4,5,6$ denotes the $\Lambda$ baryon LCDA. The mass ratio, $r=m_{\Lambda}/M_{\Lambda_{b}}$, reflects the $1/m_{b}$ power suppression associated with the high-twist $\Lambda$ baryon LCDAs. By comparing the $h_{43}$ and $h_{45}$ terms in Table 5, it is found that $h_{45}$ is power suppressed by factor of $r^{2}$ compared to $h_{43}$. However, $h_{43}$ includes an additional momentum fraction $x_{3}$.
Fig. 2 displays the dependence of the $\Lambda_{b}$ baryon LCDAs on the momentum fractions $x_{1}$, $x_{2}$ and $x_{3}$ in the exponential model. The $\Lambda_{b}$ twist-4 LCDA $\Psi_{4}$ shows a pronounced peak when $x_{1}\approx 1$ and $x_{2,3}$ approach zero, clearly indicating that the b-quark carries most of the baryon’s momentum. The presence of a small momentum fraction $x_{3}$ closing to zero significantly suppresses the contribution of $h_{43}$. Similarly, near-vanishing momentum fraction $x_{3}$ also significantly suppresses the numerical results of $h_{3^{+-}4}$ and $h_{3^{-+}4}$. These results demonstrate that high-power contributions to the baryon transition form factors receive substantial endpoint enhancement, emerging as the dominant effect, which is in agreement with the conclusions of Ref. Wang:2011uv ; Han:2022srw .
Table 5: Numerators $h_{ij}$ of the integrands in the factorization formula for the form factor $f_{1}^{V}(0)$ with $i=2,3^{+-},3^{-+},4$ and $j=3,4,5,6$. an overall coefficient $\mathcal{C}m_{\Lambda_{b}}^{3}f_{\Lambda_{b}}/(128N_{c}^{2})$ with the color factor $\mathcal{C}=8/3$ is implicit for the entries.
@@
Figure 2: Dependencies of the $\Lambda_{b}$ baryon LCDAs $\psi_{2}$, $\psi_{3}^{+-}$, $\psi_{3}^{-+}$ and $\psi_{4}$ on the momentum fractions $x_{i}$ for the exponential model in Bell:2013tfa . Each point inside the triangles satisfies the relation $x_{1}+x_{2}+x_{3}=1$.
To calculate the observables of semileptonic decays, we require the behavior of the $\Lambda_{b}\to\Lambda$ transition form factors across the whole kinematic range $0\leq q^{2}\leq q_{max}^{2}=(m_{\Lambda_{b}}-m_{\Lambda})^{2}$. The form factors are first evaluated at three $q^{2}$ points in the low-$q^{2}$ region, $0\leq q^{2}\leq m_{\tau}^{2}$, where PQCD approach is reliable, and then extended to the high-$q^{2}$ regime using the z-series formula Bourrely:2008za .
$F(q^{2})=\frac{1}{1-q^{2}/(m_{\rm pole}^{F})^{2}}\sum_{i=0}^{k}\left(a_{i}^{F}[z(q^{2})]^{i}\right).$ (35)
where $F\equiv f_{+,\perp,0},g_{+,\perp,0},h_{+,\perp},\widetilde{h}_{+,\perp}$, the parameter $z$ is defined as,
@@
The central values and uncertainties of the parameters $a_{0}^{F}$ and $a_{1}^{F}$, obtained through a global fit using the three points at low-$q^{2}$ region calculated by the PQCD approach and one point at $q_{max}^{2}$ from the lattice QCD Detmold:2016pkz , are listed in Table 7. The form factors over the full kinematic region are presented in Fig. 3.

### v2

@@ III Light-cone distribution amplitudes
Refs. Bali:2015ykx ; RQCD:2019hps systematically research the normalization constants and first moments of the leading twist distribution amplitudes of the baryon octet in the lattice QCD. Leading twist DAs can be expanded in a set of orthogonal polynomials under conformal partial wave expansion as,
$\begin{split}V_{1}(x_{1},x_{2},x_{3})&=120x_{1}x_{2}x_{3}(x_{1}-x_{2})\left(\frac{21\sqrt{6}}{4}\phi_{11}-\frac{7\sqrt{6}}{4}\phi_{10}\right)\\ A_{1}(x_{1},x_{2},x_{3})&=120x_{1}x_{2}x_{3}\left(-f_{\Lambda}+\frac{7\sqrt{6}}{4}(\phi_{11}+\phi_{10})(x_{1}+x_{2})-\frac{14\sqrt{6}}{4}x_{3}(\phi_{11}+\phi_{10})\right)\\ T_{1}(x_{1},x_{2},x_{3})&=120x_{1}x_{2}x_{3}\pi_{10}\frac{7\sqrt{6}}{2}(x_{1}-x_{2}).\end{split}$ (28)
The leading contributions $120x_{1}x_{2}x_{3}f_{\Lambda}$ is usually referred to as the asymptotic DAs with the normalization coefficient $f_{\Lambda}$. Higher-order coefficients $\phi_{10}$, $\phi_{11}$ and $\pi_{10}$ are the first moments and usually referred to as shape parameters. Those parameters, rescaled to $1GeV$, are obtained in the the continuum limit, $f_{\Lambda}=6.20^{+0.11}_{-0.09}\times 10^{-3}GeV^{2}$, $\phi_{11}=0.249^{+0.058}_{-0.058}\times 10^{-3}GeV^{2}$, $\phi_{10}=0.719^{+0.040}_{-0.044}\times 10^{-3}GeV^{2}$ and $\pi_{10}=0.252^{+0.040}_{-0.033}\times 10^{-3}GeV^{2}$ RQCD:2019hps .
For the high-twist DAs $V_{i},A_{i},T_{i},S_{i}$ and $P_{i}$ of the $\Lambda$ baryon, we employ the parameterization given in Ref. Liu:2009mb , which is directly extended from the nucleon LCDAs in Ref. Braun:2000kw under the constraints of the equations of motion and general symmetry,
$S_{1}(x_{1},x_{2},x_{3})$ $=$ $6x_{3}(1-x_{3})(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{1}(x_{1},x_{2},x_{3})=6x_{3}(1-x_{3})(\xi_{4}^{0}-\xi_{4}^{{}^{\prime}0})\,$,
$V_{2}(x_{1},x_{2},x_{3})$ $=$ $0\,,\hskip 142.26378ptA_{2}(x_{1},x_{2},x_{3})=-24x_{1}x_{2}\phi_{4}^{0}\,$,
$V_{3}(x_{1},x_{2},x_{3})$ $=$ $12(x_{1}-x_{2})x_{3}\psi_{4}^{0}\,,\hskip 65.44142ptA_{3}(x_{1},x_{2},x_{3})=-12x_{3}(1-x_{3})\psi_{4}^{0}\,$,
$T_{2}(x_{1},x_{2},x_{3})$ $=$ $0\,,\hskip 142.26378ptT_{3}(x_{1},x_{2},x_{3})=6(x_{2}-x_{1})x_{3}(-\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,$,
$T_{7}(x_{1},x_{2},x_{3})$ $=$ $-6(x_{1}-x_{2})x_{3}(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,$
$S_{2}(x_{1},x_{2},x_{3})$ $=$ $\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{2}(x_{1},x_{2},x_{3})=\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,$,
$V_{4}(x_{1},x_{2},x_{3})$ $=$ $3(x_{2}-x_{1})\psi_{5}^{0}\,,\hskip 79.6678ptA_{4}(x_{1},x_{2},x_{3})=-3(1-x_{3})\psi_{5}^{0}\,$,
$V_{5}(x_{1},x_{2},x_{3})$ $=$ $0\,,\hskip 139.41832ptA_{5}(x_{1},x_{2},x_{3})=-6x_{3}\phi_{5}^{0}\,$,
$T_{4}(x_{1},x_{2},x_{3})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 36.98866ptT_{5}(x_{1},x_{2},x_{3})=0\,$,
$T_{8}(x_{1},x_{2},x_{3})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,$
$V_{6}(x_{1},x_{2},x_{3})$ $=$ $0\,,\hskip 142.26378ptA_{6}(x_{1},x_{2},x_{3})=-2\phi_{6}^{0}\,$,
$T_{6}(x_{1},x_{2},x_{3})$ $=$ $0\,.$ (29)
The recently developed large-momentum effective theory (LaMET) enables extraction of fully $x$-dependent distribution amplitudes from first principles on the lattice LatticeParton:2024vck ; LPC:2025jvd , including the contribution of higher moments, which could provide more reliable inputs for future precision calculations of baryons.
The parameters $\phi_{i}^{0}$, $\psi_{i}^{0}$, $\xi_{i}^{0}$ and $\xi_{i}^{\prime 0}$ in the above DAs connect to the $f_{\Lambda}$ and three new P-wave (high twist) normalization constants $\lambda_{1,2,3}$ as Liu:2009mb ; Bali:2015ykx ,
$\phi_{6}^{0}$ $=$ $f_{\Lambda},\hskip 133.72786pt\phi_{4}^{0}=\phi_{5}^{0}=-\frac{1}{2}(f_{\Lambda}-\frac{1}{\sqrt{6}}\lambda_{1}),$
$\psi_{4}^{0}$ $=$ $\psi_{5}^{0}=\frac{1}{2}(f_{\Lambda}+\frac{1}{\sqrt{6}}\lambda_{1}),\hskip 45.52458pt\xi_{4}^{0}=\xi_{5}^{0}=-\frac{\lambda_{2}}{2\sqrt{6}},$
$\xi_{4}^{{}^{\prime}0}$ $=$ $\xi_{5}^{{}^{\prime}0}=-\frac{\lambda_{3}}{\sqrt{6}}.$ (30)
$\lambda_{1,2,3}$ can be defined as matrix elements of local three-quark twist-four operators without derivatives Braun:2000kw ; Wein:2015oqa ,
$\begin{split}\langle 0|\left(u^{T}(0)C\gamma^{\mu}\gamma_{5}d(0)\right)\gamma_{\mu}s(0)|\Lambda(p,\lambda)\rangle=&\frac{-1}{\sqrt{6}}\lambda_{1}m_{\Lambda}u_{\Lambda}(p,\lambda),\\ \langle 0|\left(u^{T}(0)Cd(0)\right)\gamma_{5}s(0)|\Lambda(p,\lambda)\rangle=&\frac{-1}{4\sqrt{6}}(\lambda_{2}+2\lambda_{3})m_{\Lambda}u_{\Lambda}(p,\lambda),\\ \langle 0|\left(u^{T}(0)C\gamma_{5}d(0)\right)s(0)|\Lambda(p,\lambda)\rangle=&\frac{-1}{4\sqrt{6}}(\lambda_{2}-2\lambda_{3})m_{\Lambda}u_{\Lambda}(p,\lambda).\end{split}$ (31)
The explicit values from lattice QCD calculation for the parameters $\lambda_{1,2,3}$ are $\lambda_{1}=(-37.6\pm 3.5)\times 10^{-3}GeV^{2}$, $\lambda_{2}=(88.3\pm 4.6)\times 10^{-3}GeV^{2}$ and $\lambda_{3}=(-46.7\pm 4.5)\times 10^{-3}GeV^{2}$ RQCD:2019hps .
@@
The PQCD factorization formulas for all relevant Feynman diagrams are provided in Appendix Appendix: Factorization formulas. We present our numerical results for the $\Lambda_{b}\to\Lambda$ form factors at $q^{2}=0\>\text{GeV}^{2}$ in Table 1, and compare with those from the lattice QCD Detmold:2016pkz . The uncertainties in this work arise from the parameters describing the $\Lambda_{b}$ and $\Lambda$ baryon LCDAs. Further improvements in the determination of baryon LCDAs will help reduce theoretical uncertainties in future predictions.
As shown in Table 1, our results are consistent with those from the lattice QCD within uncertainties and satisfy the expected endpoint relations $f_{0}(0)=f_{+}(0)$ and $g_{0}(0)=g_{+}(0)$. For further comparison with the form factors from other approaches, we convert the obtained helicity form factors in Table 1 to Weinberg form factors in Table 2. Our predictions are consistent with expectations from heavy-quark effective theory (HQET), which predicts for heavy-to-light baryon transitions form factors at $q^{2}=0$, $f_{1}^{V}(0)\approx f_{1}^{A}(0)\gg f_{2,3}^{V,A}(0)$, $f_{2}^{TV}(0)\approx f_{2}^{TA}(0)\gg f_{1}^{TV,TA}(0)$.
Table 1: Form factors of the $\Lambda_{b}\to\Lambda$ transition obtained in the PQCD and lattice QCD in helicity basis at $q^{2}=0GeV^{2}$.
this work lattice QCD Detmold:2016pkz
$f_{+}(0)$ $0.149\pm 0.035$ $0.160\pm 0.024$
$f_{\perp}(0)$ $0.185\pm 0.043$ $0.207\pm 0.037$
$f_{0}(0)$ $0.149\pm 0.036$ $0.156\pm 0.035$
$g_{+}(0)$ $0.159\pm 0.038$ $0.112\pm 0.028$
$g_{\perp}(0)$ $139\pm 0.034 0.0$$095\pm 0.035$
$g_{0}(0)$ $0.159\pm 0.038$ $0.166\pm 0.023$
$h_{+}(0)$ $0.185\pm 0.043$ $0.236\pm 0.041$
$h_{\perp}(0)$ $0.154\pm 0.037$ $0.166\pm 0.022$
$\widetilde{h}_{+}(0)$ $0.138\pm 0.034$ $0.163\pm 0.029$
$\widetilde{h}_{\perp}(0)$ $0.154\pm 0.036$ $0.156\pm 0.027$
Table 2: Theoretical predictions for the $\Lambda_{b}\rightarrow\Lambda$ form factors in Weinberg basis at $q^{2}=0$. The symbol $\cdots$ denotes that there is no relevant result in literature.
@@
this work $149\pm 0.036 0.0$$30\pm 0.006 0.0$$015\pm 0.003$ $159\pm 0.039 0.0$$25\pm 0.006 0.0$$020\pm 0.004$
Lattice QCDDetmold:2016pkz $158\pm 0.004 0.0$$41\pm 0.005 0.0$$017\pm 0.009$ $137\pm 0.004 0.0$$055\pm 0.008$ $0.117\pm 0.009$
@@
this work $0.015\pm 0.006$ $154\pm 0.037 0.0$$019\pm 0.004$ $0.154\pm 0.036$
Lattice QCDDetmold:2016pkz $0.059\pm 0.006$ $166\pm 0.005 0.0$$009\pm 0.008$ $0.156\pm 0.005$
@@
Contributions to the form factors $f_{1}^{V,A}(0)$ from each Feynman diagram are presented in Table 3, which reveals that the diagrams $d_{7}$ and $d_{8}$ dominate. We analyze the contributions of each twist-term combination for these two diagrams in Table 4, which shows that the $\Lambda_{b}$ twist-4 and $\Lambda$ twist-5 combination accounts for the dominant numerical results in both $d_{7}$ and $d_{8}$. These results clearly demonstrate the essential role of higher-power effects in baryon decays at the $b$-quark mass scale.
@@
$f_{1}^{V}(0)$ $4.7\times 10^{-4}$ $4.6\times 10^{-4}$ $6.5\times 10^{-3}$ $7.1\times 10^{-4}$ $3.5\times 10^{-3}$ $4.2\times 10^{-3}$ 0.107 0.093
$f_{1}^{A}(0)$ $3.4\times 10^{-4}$ $2.1\times 10^{-5}$ $5.9\times 10^{-3}$ $5.7\times 10^{-4}$ $3.5\times 10^{-3}$ $3.6\times 10^{-3}$ 0.106 0.092
@@
$f_{1}^{V}(0)$ $5.8\times 10^{-3}$ $4.6\times 10^{-4}$ 0.021 $1.1\times 10^{-3} -$0.028 -0.034 0.149
$f_{1}^{A}(0)$ $5.9\times 10^{-3}$ $3.8\times 10^{-4}$ 0.021 $1.1\times 10^{-3} -$0.024 -0.029 0.159
@@
$\Lambda_{b}$ twist-4 -0.017 $\sim 0$ 0.118 $\sim 0$ -0.015 $\sim 0$ 0.103 $\sim 0$
Recalling that the form factors are expressed as the convolution of the perturbatively calculated hard kernels and the non-perturbative baryon LCDAs. We list the numerators $h_{ij}$ of the hard kernels for the diagram $d_{7}$ contributing to the form factor $f_{1}^{V}(0)$ in Table 5, classified by different twist combination. $i=2,3,4$ stands for the $\Lambda_{b}$ baryon LCDA and $j=3,4,5,6$ denotes the $\Lambda$ baryon LCDA. The mass ratio, $r=m_{\Lambda}/M_{\Lambda_{b}}$, reflects the $1/m_{b}$ power suppression associated with the high-twist $\Lambda$ baryon LCDAs. By comparing the $h_{43}$ and $h_{45}$ terms in Table 5, it is found that $h_{45}$ is power suppressed by factor of $r^{2}$ compared to $h_{43}$. However, $h_{43}$ includes an additional momentum fraction $x_{3}$.
Fig. 2 displays the dependence of the $\Lambda_{b}$ baryon LCDAs on the momentum fractions $x_{1}$, $x_{2}$ and $x_{3}$ in the exponential model. The twist-4 $\Lambda_{b}$ LCDA $\Psi_{4}$ shows a marked peak where $x_{1}\approx 1$ and $x_{2,3}$ approach zero, clearly indicating that the b-quark carries most of the baryon’s momentum. The presence of the momentum fraction $x_{3}$ in $h_{43}$ closing to zero significantly suppresses the contribution from the endpoint peak region of the twist-4 $\Lambda_{b}$ LCDA $\Psi_{4}$. This pattern demonstrates that high-power contributions to the baryon transition form factors are significantly enhanced near the endpoint region, emerging as the dominant effects, which is in agreement with the earlier findings in Ref. Wang:2011uv ; Han:2022srw .
Table 5: Numerators $h_{ij}$ of the hard kernels in the factorization formula for the form factor $f_{1}^{V}(0)$ with $i=2,3^{+-},3^{-+},4$ and $j=3,4,5,6$. An overall coefficient $\mathcal{C}m_{\Lambda_{b}}^{3}f_{\Lambda_{b}}/(128N_{c}^{2})$ with the color factor $\mathcal{C}=8/3$ is implicit for each entry.
@@
Figure 2: Dependencies of the $\Lambda_{b}$ baryon LCDAs $\psi_{2}$, $\psi_{3}^{+-}$, $\psi_{3}^{-+}$ and $\psi_{4}$ on the momentum fractions $x_{i}$ in the exponential model Bell:2013tfa . Each point inside the triangles satisfies the relation $x_{1}+x_{2}+x_{3}=1$.
To calculate the observables of semileptonic decays, extrapolation of the $\Lambda_{b}\to\Lambda$ form factors are performed to the whole kinematic range $0\leq q^{2}\leq q_{max}^{2}=(m_{\Lambda_{b}}-m_{\Lambda})^{2}$. The form factors are firstly evaluated at six $q^{2}$ points in the low-$q^{2}$ region, $0\leq q^{2}\leq m_{\tau}^{2}$, where the PQCD approach remains reliable, and then extrapolated to high-$q^{2}$ regime using the z-series formula Bourrely:2008za ,
$F(q^{2})=\frac{1}{1-q^{2}/(m_{\rm pole}^{F})^{2}}\sum_{i=0}^{k}\left(a_{i}^{F}[z(q^{2})]^{i}\right),$ (32)
where $F=f_{+,\perp,0},g_{+,\perp,0},h_{+,\perp},\widetilde{h}_{+,\perp}$, and the parameter $z$ is defined as,
@@
To constrain the behavior at high $q^{2}$ region, the form factors at $q^{2}_{max}$ from the lattice QCD calculation Detmold:2016pkz are included into our global fit.
@@
The central values and uncertainties of the fit parameters $a_{0}^{F}$ and $a_{1}^{F}$ are listed in Table 7. The corresponding form factors at the full kinematic region are depicted in Fig. 3.